Ravi's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs $5.45 Ravi per pound, and type B coffee costs $4.20 per pound. This month, Ravi made 121 pounds of the blend, for a total cost of $571.95 . How many pounds of type B coffee did he use?
Answer by mananth(16946)

x= pounds of coffee type I
y= pounds of coffee type II

1.00 x + 1.00 y = 121.00 .............1
Total value
5.45 x + 4.20 y = 571.95 .............2
Eliminate y
multiply (1)by -4.20
Multiply (2) by 1.00
-4.20 x -4.20 y = -508.20
5.45 x + 4.20 y = 571.95
Add the two equations
1.25 x = 63.75
/ 1.25
x = 51.00
plug value of x in (1)
1.00 x + 1.00 y = 121.00
51.00 + y = 121.00
y = 121.00 -51.00
y = 70.00
y = 70.00
x= 51.00 pounds of coffee type I
y= 70.00 pounds of coffee type II